A lift is ascending with acceleration g/3. What will be the time period of a simple pendulum suspended from its ceiling if its time period in stationary lift is

Grade:12th pass

2 Answers

Nitin Bindal
7 years ago
If its tp im sationaray lift is T and T=2πrootl/g but in a moving lift it becomesT=2πroot(l/g+g/3). NOW SOLVE
Sanjeev Gupta
6 years ago
g+g/3 = 4g/3- equation 1. now T=2pie root l/g eqution 2Puting value from equation 1T=2pie root l/4g/3 Now T=2pie root 3l/ 4g- equation 3Now equating eq. 2and 3 t=root 3/4 T is the ans